### v1

# Curvatures and Non-metricities in the Non-Relativistic Limit of Bosonic Supergravity

Abstract
We construct a diffeomorphism-covariant formulation of the non-relativistic (NR) limit of bosonic supergravity. This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to the NR degrees of freedom. The construction is purely geometrical and is based on a torsionless connection. The non-metricities are associated with the gravitational fields of the theory, $\tau_{\mu\nu},h_{\mu\nu},\tau^{\mu\nu}$ and $h^{\mu\nu}$, and are fixed by requiring compatibility with the relativistic metric. We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature. Our results establish an equivalence between the proposed construction and the intrinsic torsion framework of string Newton–Cartan geometry. We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
I Introduction
In recent years, non-relativistic (NR) limits of string theory NR1 -NR3 and supergravity NRST1 -NRST13 have attracted considerable attention, motivated by both conceptual and more phenomenological developments. In particular, the systematic understanding of the supergravity limit of NR string theory has led to the emergence of Newton–Cartan and string Newton–Cartan geometries as the natural geometric frameworks underlying these limits (for reviews and complementary introductions, see Review -Review3 and references therein). Let us start by considering the universal bosonic NS–NS supergravity, whose relativistic field content consists of a spacetime metric $\hat{g}_{\mu\nu}$, a Kalb–Ramond two-form $\hat{B}_{\mu\nu}$ and a dilaton $\hat{\phi}$. A controlled NR limit is obtained NSNS by introducing an explicit expansion in powers of the speed of light $c$,
$\hat{g}_{\mu\nu}$ $=$ $c^{2}\tau_{\mu\nu}+h_{\mu\nu}\,$, (1)
$\hat{g}^{\mu\nu}$ $=$ $\frac{1}{c^{2}}\tau^{\mu\nu}+h^{\mu\nu}\,$, (2)
$\hat{B}_{\mu\nu}$ $=$ $-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,$, (3)
$\hat{\phi}$ $=$ $\ln(c)+\varphi\,$, (4)
where $\mu,\nu=0,\dots,25$ are space-time indices and the longitudinal directions are labeled by $a,b=0,1$. We also have
$\tau_{\mu\nu}$ $=$ $\tau_{\mu}{}^{a}\tau_{\nu}{}^{b}\eta_{ab}\,$, (5)
$\tau^{\mu\nu}$ $=$ $\tau^{\mu}{}_{a}\tau^{\nu}{}_{b}\eta_{ab}\,$, (6)
$h_{\mu\nu}$ $=$ $e_{\mu}{}^{a^{\prime}}e_{\nu}{}^{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (7)
$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (8)
according to the vielbein decomposition of the fields ($a^{\prime}=2,...,25$). Moreover, they satisfy the defining relations,
$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,$,
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
Although individual geometric quantities constructed from the relativistic Levi–Civita connection exhibit divergences in the limit $c\to\infty$, the bosonic NS–NS action,
$S=\int d^{26}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big)\,$, (10)
with $\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}$, remains finite due to a non-trivial cancellation between divergent contributions arising from the Ricci scalar and the Kalb–Ramond sector. This observation has been central to the development of other NR string supergravity theories such as heterotic formulations BandR - Sigmaheterotic , supersymmetric extensions NRST11 -NRST12 and Double Field Theory DFT1 -DFT4 , in the context of non-Riemannian double geometries NRDFT1 -NRDFT9 and its NR limit EandD .
Motivated by the progress on the NR limit of string supergravities, different approaches to study higher-derivative corrections under the NR limit have recently emerged Higher-Derivative1 -Higher-Derivative4 . Some of the theories already explored can be written in terms of tensors with respect to infinitesimal diffeomorphisms in their relativistic level. In order to be able to decompose these tensors preserving the covariance, one needs to construct a suitable connection with respect to the longitudinal and transverse parts of the metric degrees of freedom (and its inverse), and rewrite these contributions in terms of curvatures and covariant derivatives with respect to this symmetry. The same can be done for the Lorentz inherited symmetries, like boosts, $SO(1,1)$ and $SO(8)$ transformations, which was done for the NR limit of bosonic supergravity in NSNS at the two-derivative level. In particular, a fully diffeomorphism-covariant formulation of this theory, formulated directly in terms of intrinsic NR geometric data and their associated curvatures is still lacking. The main obstruction lies in the fact that the relativistic Levi–Civita connection does not admit a smooth NR limit when expressed solely in terms of Newton–Cartan variables.
In this work, we solve this issue by constructing a torsionless affine connection that is explicitly covariant under spacetime diffeomorphisms. Crucially, this connection is not metric compatible with respect to the fundamental NR fields. Instead, the covariant derivatives of $\tau_{\mu}{}^{a}$ and $h^{\mu\nu}$ are controlled by specific non-metricity tensors, which are uniquely fixed by requiring consistency with the relativistic metric condition. As a result, the full bosonic supergravity action can be rewritten in a manifestly covariant form with respect to diffeomorphisms, where all terms remain finite and well defined as $c\rightarrow\infty$.
The purpose of this paper is threefold. First, we present an explicit construction of a torsionless connection and the associated non-metricities within the framework of Newton–Cartan gravity. Second, we compute the corresponding curvatures and show explicitly how all relativistic curvature tensors reorganize into NR geometric invariants. Third, we discuss several applications of the formalism. These include the derivation of the complete non-relativistic bosonic supergravity Lagrangian in a fully covariant form, the systematic analysis of higher-derivative contributions such as Riemann-squared and other four-derivative bosonic corrections, and the construction of more general Newton–Cartan–inspired theories with alternative choices of non-metricity that relax boost invariance beyond the bosonic supergravity setting.
Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity. We expect this framework to serve as a useful starting point for higher-derivative extensions and for exploring duality-covariant formulations of string theory in the NR limit.
II Infinitesimal diffeomorphism transformations and covariant derivatives
We start by recalling the transformation rule of the fundamental fields with respect to infinitesimal diffeomorphisms,
$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau_{\mu}{}^{a}}+\partial_{\mu}{\xi^{\rho}}\tau_{\rho}{}^{a}\,$,
$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau^{\mu}{}_{a}}-\partial_{\rho}{\xi^{\mu}}\tau^{\rho}{}_{a}\,$,
$\delta_{\xi}h_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h_{\mu\nu}}+2\partial_{(\mu}{\xi^{\rho}}h_{\nu)\rho}\,$,
$\delta_{\xi}h^{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h^{\mu\nu}}–2\partial_{\rho}{\xi^{(\mu}}h^{\nu)\rho}\,.$ (11)
The torsionless connection is given by
$\Gamma_{\mu\nu}^{\rho}(\tau,h)=\frac{1}{2}h^{\rho\sigma}(2\partial_{(\mu}h_{\nu)\sigma}-\partial_{\sigma}h_{\mu\nu})$
[…]
which is the $c^{0}$-contribution coming from the relativistic connection, preserving the non-tensorial transformation rule, in the sense that it is the contribution which transforms as a connection.
The covariant derivatives are therefore defined as
$\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}\,$, (13)
$\nabla_{\mu}\tau^{\nu}{}_{a}$ $=$ $\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\sigma}^{\nu}\tau^{\sigma}{}_{a}\,$, (14)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $\partial_{\mu}h_{\nu\rho}-2\Gamma_{\mu(\nu}^{\sigma}h_{\rho)\sigma}\,$, (15)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $\partial_{\mu}h^{\nu\rho}+2\Gamma_{\mu\sigma}^{(\nu}h^{\rho)\sigma}\,.$ (16)
The Riemann tensor can be easily computed from the commutator of the covariant derivatives,
$[\nabla_{\mu},\nabla_{\nu}]\tau^{\rho}{}_{a}=R^{\rho}{}_{\epsilon\mu\nu}\tau^{\epsilon}{}_{a}\,$, (17)
giving the usual expression
$R^{\rho}{}_{\epsilon\mu\nu}(\tau,h)=2\partial_{[\mu}\Gamma_{\nu]\epsilon}^{\rho}+2\Gamma_{[\mu|\alpha}^{\rho}\Gamma_{|\nu]\epsilon}^{\alpha}\,.$ (18)
III Non-metricities
The non-metricities play a fundamental role in the theory, since they guarantee that the NR limit can be taken from the relativistic theory,
$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $Q^{(\tau)}_{\mu\nu\rho}\,$, (19)
$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $Q^{(\tau^{-1})}_{\mu}{}^{\nu\rho}\,$, (20)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $Q^{(h)}_{\mu\nu\rho}\,$, (21)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $Q^{h^{-1}}_{\mu}{}^{\nu\rho}\,$. (22)
The fixed values of the non-metricities come from the metric conditions,
$\hat{\nabla}_{\mu}g_{\nu\rho}$ $=$ $0\,$,
$\hat{\nabla}_{\mu}g^{\nu\rho}$ $=$ $0\,$,
where $\hat{\nabla}$ is the usual covariant derivative with respect the relativistic Levi–Civita connection. Therefore, after doing the NR expansion, one needs the following non-metricity combinations,
$Q^{(\tau)}_{\mu\nu\rho}$ $=$ $e^{\alpha a^{\prime}}(\partial_{\mu}\tau_{(\nu\alpha}+\partial_{(\nu}\tau_{\mu\alpha}-\partial_{\alpha}\tau_{\mu(\nu})e_{\rho)a^{\prime}}$ (23)
$Q^{(\tau^{-1})}_{\rho}{}^{\mu\nu}$ $=$ $-\tau^{(\mu\alpha}(2\partial_{(\rho}h_{\sigma)\alpha}-\partial_{\alpha}h_{\rho\sigma})h^{\sigma\nu)}$ (24)
$Q^{(h)}_{\mu\nu\rho}$ $=$ $\tau^{\alpha}{}_{b}(\partial_{\mu}h_{(\nu\alpha}+\partial_{(\nu}h_{\mu\alpha}-\partial_{\alpha}h_{\mu(\nu})\tau_{\rho)}{}^{b}$ (25)
$Q^{h^{-1}}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
The previous quantities are not boost invariants  [^1] Here $\lambda_{aa^{\prime}}$ is a generic boost parameter NSNS .,
$\delta_{\lambda}h_{\mu\nu}$ $=$ $-2\lambda_{aa^{\prime}}e_{(\mu}{}^{a^{\prime}}\tau_{\nu)}{}^{a}\,$, (27)
$\delta_{\lambda}\tau^{\mu\nu}$ $=$ $2\lambda_{a}{}^{a^{\prime}}e^{(\mu}{}_{a^{\prime}}\tau^{\nu)a}\,$, (28)
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-2\epsilon_{ab}\lambda^{a}{}_{a^{\prime}}\tau_{[\mu}{}^{b}e_{\nu]}{}^{a^{\prime}}\,$. (29)
and consequently the non-metricities cannot be set to zero without compromising the boost symmetry in bosonic supergravity. We will return to this point in the discussion section, since one might be interested in exploring Newton-Cartan geometries with more arbitrary non-metricities, beyond bosonic supergravity.
IV Decomposition of the relativistic curvatures
IV.1 The Riemann tensor
The relativistic Riemann tensor is given by
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{(4)}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{(2)}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
$+c^{(-2)}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{(-4)}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
So far we have just identified part of the covariant contribution inside $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$, given by $R^{\rho}{}_{\epsilon\mu\nu}$ but other covariant contributions are encoded in $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ due to $c^{2}$- and $c^{-2}$-contributions in the relativistic Levi-Civita connection. All the contributions in the relativistic Riemann tensor can be written in covariant form, and this is what we are going to show now. Let us start with the higher-order contribution,
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\nu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
$(\partial_{\nu]}\tau_{\epsilon\beta}+\partial_{\epsilon}\tau_{\nu]\beta}-\partial_{\beta}\tau_{\nu]\epsilon})\,.$
The previous quantity transforms as a tensor, and it can be written in a manifest covariant form using the covariant derivatives,
$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
$(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
The same technique can be implemented for the lowest contribution of the Riemann tensor,
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\nu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
$(\partial_{\nu]}h_{\epsilon\beta}+\partial_{\epsilon}h_{\nu]\beta}-\partial_{\beta}h_{\nu]\epsilon})$
which therefore can be written as
$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
$(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$
In these cases we have identified that the structure of $R^{(-4)\rho}{}_{\epsilon\mu\nu}$ is just an interchange of $\tau_{\mu\nu}\leftrightarrow h_{\mu\nu}$ and $h^{\mu\nu}\leftrightarrow\tau^{\mu\nu}$ with respect to the highest order contribution, $R^{(4)\rho}{}_{\epsilon\mu\nu}$.
The $c^{2}$-contributions of the relativistic Riemann tensor takes an important role, since it is part of bosonic supergravity action once one takes the NR limit. These contributions can be written in terms of the covariant derivatives
$\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}=2\nabla_{[\mu}\Big[h^{\rho\alpha}(\nabla_{\nu]}\tau_{\epsilon\alpha}+\nabla_{\epsilon}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\epsilon})\Big]\,.$ (35)
Again, we can interchange $\tau_{\mu\nu}\leftrightarrow h_{\mu\nu}$ and $h^{\mu\nu}\leftrightarrow\tau^{\mu\nu}$ to construct the $c^{(-2)}$-contributions,
$\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}=2\nabla_{[\mu}\Big[\tau^{\rho\alpha}(\nabla_{\nu]}h_{\epsilon\alpha}+\nabla_{\epsilon}h_{\nu]\alpha}-\nabla_{\alpha}h_{\nu]\epsilon})\Big]\,.$ (36)
Finally, we proceed to construct the $c^{0}$-contributions to the relativistic Riemann tensor. This quantity is given by,
$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
In the next section we will study the decomposition of the (relativistic) Ricci tensor in terms of the Newton-Cartan fields, and in covariant form with respect to infinitesimal diffeomorphisms.
IV.2 The Ricci tensor
The relativistic Ricci tensor is given by
$\hat{R}_{\epsilon\nu}$ $=$ $c^{4}\hat{R}^{(4)}_{\epsilon\nu}+c^{2}\hat{R}^{(2)}_{\epsilon\nu}+\hat{R}^{(0)}_{\epsilon\nu}+c^{-2}\hat{R}^{(-2)}_{\epsilon\nu}+c^{-4}\hat{R}^{(-4)}_{\epsilon\nu}\,$. (38)
Using the results of the previous subsection, we can directly write all the contributions in covariant form,
$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
$\hat{R}^{(2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[h^{\mu\alpha}(\nabla_{\nu]}\tau_{\epsilon\alpha}+\nabla_{\epsilon}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\epsilon})\Big]\,$, (40)
$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
$\hat{R}^{(-2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[\tau^{\mu\alpha}(\nabla_{\nu]}h_{\epsilon\alpha}+\nabla_{\epsilon}h_{\nu]\alpha}-\nabla_{\alpha}h_{\nu]\epsilon})\Big]\,.$ (42)
$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
IV.3 The Ricci scalar
The relativistic Ricci scalar is decomposed as
$\hat{R}$ $=$ $c^{2}\hat{R}^{(2)}+\hat{R}^{(0)}+c^{-2}\hat{R}^{(-2)}+c^{-4}\hat{R}^{(-4)}\,$, (44)
where the covariant form of each contribution is given by
$\hat{R}^{(2)}$ $=$ $\hat{R}^{(2)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(4)}_{\mu\nu}\tau^{\mu\nu}\,$, (45)
$\hat{R}^{(0)}$ $=$ $\hat{R}^{(0)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(2)}_{\mu\nu}\tau^{\mu\nu}\,$, (46)
$\hat{R}^{(-2)}$ $=$ $\hat{R}^{(-2)}_{\mu\nu}h^{\mu\nu}+\hat{R}^{(0)}_{\mu\nu}\tau^{\mu\nu}\,$, (47)
$\hat{R}^{(-4)}$ $=$ $\hat{R}^{(-4)}_{\mu\nu}h^{\mu\nu}\,.$ (48)
Using the results of the previous subsection we can easily see that the right hand side of the previous contributions can be written in a fully covariant way. We will explore the explicit form of $\hat{R}^{(0)}$ in the next section.
V Applications
In this section, we explore three direct applications of the proposed formalism. First, we derive the finite non-relativistic limit of the two-derivative bosonic supergravity Lagrangian. We then discuss the convenience of this formulation for extracting covariant contributions at any desired order in the expansion in powers of $c$. As an illustrative example, we compute all finite contributions arising from the Metsaev–Tseytlin Lagrangian, which constitutes part of the four-derivative corrections to the bosonic Lagrangian in the NR limit [^2] Additional contributions are expected to arise in order to cancel the higher-derivative divergences of the action.. Finally, we discuss the possibility of employing more general non-metricities, focusing in particular on the case in which all such tensors vanish. This choice explicitly breaks the boost symmetry of the NR limit of bosonic supergravity.
V.1 NR-Bosonic supergravity Lagrangian: covariant (re-)formulation
This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the $\hat{H}-$terms, and using the results of the previous section it can be written in the following way,
$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
$+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}$
$+2\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\nu]}\tau_{\mu\alpha}+\nabla_{\mu}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi$
$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
where
$h_{\mu\nu\rho}$ $=$ $3\partial_{[\mu}b_{\nu\rho]}\,,\quad\hat{H}^{(2)}_{\mu\nu\rho}=-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,$.
Now we can explicitly see that the finite bosonic Lagrangian after taking the NR limit contains a $R_{\mu\nu}h^{\mu\nu}$ contribution, plus other terms related to the non-metricities. The construction is equivalent to the torsionful-formalism of string Newton-Cartan.
V.2 Non-relativistic bosonic $\alpha^{\prime}$-corrections
The four-derivative corrections to the bosonic string supergravity were historically computed considering three- and four-point scattering amplitudes for the massless states GrossSloan -CN (see Electure for a pedagogical review of this topic). The method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
$S_{MT}=\int d^{D}x\sqrt{-g}e^{-2\phi}(L^{(0)}+L^{(1)}_{MT})\,$, (50)
where
$L^{(1)}_{\rm MT}$ $=$ $-\frac{\alpha^{\prime}}{4}\Big[\hat{R}_{\mu\nu\rho\sigma}\hat{R}^{\mu\nu\rho\sigma}-\frac{1}{2}\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}+\frac{1}{24}\hat{H}^{4}-\frac{1}{8}\hat{H}^{2}_{\mu\nu}\hat{H}^{2\mu\nu}\Big]\,$,\, (51)
and
$\hat{H}^{2}_{\mu\nu}$ $=$ $\hat{H}_{\mu}{}^{\rho\sigma}\hat{H}_{\nu\rho\sigma}\,,\quad\hat{H}^{2}=\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\,.$ (52)
The formalism developed in this letter allows to easily inspect the NR contributions of the previous quantities without losing covariance. For example, recalling that $\alpha^{\prime}\rightarrow\frac{\alpha^{\prime}_{NR}}{c^{2}}$ Higher-Derivative3 -Higher-Derivative4 , the finite contributions coming from $\hat{R}_{\mu\nu\rho\sigma}\hat{R}^{\mu\nu\rho\sigma}$ are given by the $c^{2}$-contributions of the previous term, after the NR decomposition,
$2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}h_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+4\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(-2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(0)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}h_{\mu\gamma}h^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}h^{\rho\delta}h^{\sigma\lambda}$
$+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}h_{\mu\gamma}h^{\nu\epsilon}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}h_{\mu\gamma}h^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}\tau^{\sigma\lambda}h^{\nu\epsilon}$
$+4\hat{R}^{(0)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}h_{\mu\gamma}h^{\nu\epsilon}h^{\sigma\lambda}+\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}h_{\mu\gamma}h^{\rho\delta}h^{\sigma\lambda}$
$+\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}\tau^{\sigma\lambda}h^{\nu\epsilon}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}h^{\nu\epsilon}$
$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}$ (53)
Similarly, one can inspect the rest of the terms. For $\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}$ the finite contributions are given by
$\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\gamma\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\sigma\alpha}h_{\mu\gamma\delta}h_{\epsilon\lambda\alpha}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\rho\gamma}h^{\nu\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\epsilon\lambda}h_{\gamma\delta\alpha}$
$+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\sigma}{}_{\gamma\epsilon\delta}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}h^{\nu\epsilon}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\gamma\lambda}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\lambda}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\gamma\lambda}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\mu}{}_{\sigma\gamma\epsilon}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}\tau^{\gamma\lambda}h^{\rho\sigma}h^{\epsilon\alpha}h_{\delta\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}\tau^{\epsilon\lambda}h^{\nu\sigma}h^{\rho\alpha}h_{\delta\lambda\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\gamma\delta}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(-2)\mu}{}_{\delta\lambda\alpha}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}h^{\rho\delta}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}h^{\rho\gamma}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}\tau^{\epsilon\alpha}h^{\rho\delta}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}\tau^{\gamma\lambda}h^{\epsilon\alpha}$
$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\rho\epsilon}\,.$ (54)
The terms only containing $\hat{H}-$contributions can also be easily written in covariant form. For example, $H^{4}$ contributes with the following finite terms,
$4\hat{H}^{(2)}_{\mu\nu\rho}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\sigma\gamma\epsilon}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+6\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$+6\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+24\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\lambda\alpha}h_{\beta\chi\psi}$
$+12\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\delta\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+36\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\gamma}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$+12\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\nu\chi}h^{\rho\psi}h^{\sigma\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\beta\chi\psi}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\rho\epsilon}h^{\delta\beta}h^{\lambda\chi}h^{\alpha\psi}$
$+18\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\delta\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}\,$, (55)
while $H^{2}_{\mu\nu}H^{2\mu\nu}$ contributes with,
$4\hat{H}^{(2)}_{\mu\nu\rho}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\delta\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\sigma\gamma\delta}h_{\epsilon\alpha\chi}h_{\lambda\beta\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\sigma}h^{\rho\gamma}h^{\epsilon\lambda}h^{\alpha\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\alpha\chi}h_{\lambda\beta\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\lambda\chi}h_{\alpha\beta\psi}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\beta}h^{\chi\psi}h_{\delta\lambda\chi}h_{\alpha\beta\psi}$
$+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\delta\lambda}h^{\mu\sigma}h^{\nu\alpha}h^{\rho\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\alpha\beta}h_{\lambda\chi\psi}+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\lambda\beta}h_{\alpha\chi\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\delta}h^{\nu\lambda}h^{\rho\alpha}h^{\sigma\beta}h^{\gamma\chi}h^{\epsilon\psi}h_{\delta\beta\chi}h_{\lambda\alpha\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\delta}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\delta}h^{\rho\gamma}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+16\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\gamma}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}h_{\beta\chi\psi}$
$+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\beta}h^{\rho\chi}h^{\gamma\delta}h^{\epsilon\lambda}h^{\alpha\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\nu\chi}h^{\rho\sigma}h^{\gamma\delta}h^{\epsilon\lambda}h^{\alpha\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\nu\beta}h^{\rho\gamma}h^{\epsilon\delta}h^{\lambda\chi}h^{\alpha\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\sigma}\tau^{\delta\beta}h^{\nu\lambda}h^{\rho\alpha}h^{\gamma\chi}h^{\epsilon\psi}h_{\beta\chi\psi}$
$-8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\tau^{\mu\beta}\tau^{\sigma\chi}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h^{\epsilon\psi}h_{\beta\chi\psi}+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\rho\delta}h^{\epsilon\beta}h^{\lambda\chi}h^{\alpha\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\delta\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}$
$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}$ (56)
At the present moment we cannot control all the divergences of the full Metsaev-Tseytlin Lagrangian and for that reason we are not giving an exhaustive analysis of the four-derivative contributions of this Lagrangian in its NR limit. However, the covariant framework developed in this work could be very useful for attacking this open problem.
V.3 General f(R,Q) non-relativistic theories
While in this work we have shown that the non-metricities (23)-(26) need to be fixed in a very particular way to provide the relativistic metric conditions (III)-(III), one can construct more general NR theories with arbitrary non-metricities. One interesting example is to demand the compatibility condition on the Newton-Cartan fields,
$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $0\,$, (57)
$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $0\,$, (58)
$\nabla_{\mu}h_{\nu\rho}$ $=$ $0\,$, (59)
$\nabla_{\mu}h^{\nu\rho}$ $=$ $0\,$, (60)
which automatically implies that both the relativistic Riemann and Ricci tensor are now finite,
$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $R^{\rho}{}_{\epsilon\mu\nu}\,$, (61)
$\hat{R}_{\epsilon\nu}$ $=$ $R_{\epsilon\nu}\,$, (62)
and the more general (gravitational) Lagrangian in this setup is given by
$L_{NR}|_{Q=0}=a_{1}R_{\mu\nu}h^{\mu\nu}+a_{2}R_{\mu\nu}\tau^{\mu\nu}\,$, (63)
with $a_{1}$ and $a_{2}$ two arbitrary coefficients. In this case, the $SO(8)$ and $SO(1,1)$ symmetry from the vielbein formalism can be preserved, but the boost symmetry is broken. Further theories with these properties can be explored by changing the non-metricity contributions and including $L(R,Q)$ in the $L_{NR}$ Lagrangian.
VI Discussion
The present construction provides a geometrical formulation of the non-relativistic limit of bosonic supergravity in terms of a Newton–Cartan geometry with non-metricities. Within this framework, an appropriate choice of connection allows spacetime diffeomorphism covariance to remain manifest throughout the construction. As a result, the full two-derivative bosonic supergravity in its NR limit can be written in a covariant form with respect to infinitesimal diffeomorphisms, thereby complementing other approaches based on subgroups of the Lorentz symmetry NSNS .
A key advantage of this formalism is that it allows for a straightforward decomposition of bosonic higher-derivative contributions, such as $\alpha^{\prime}$-corrections. These corrections are naturally formulated in the metric formalism, where the Lorentz symmetry is manifest and no vielbein (or spin connection) is required. From a practical perspective, this provides a systematic method for analyzing higher-derivative terms by expressing relativistic curvature invariants directly in terms of non-relativistic geometric data. In particular, the use of fixed non-metricities enables a transparent identification of divergent and finite contributions, a task that is typically challenging.
Although we do not yet have full control over the divergences that appear at the four-derivative level, the formalism developed in this work already proves useful for isolating and organizing several finite contributions to the four-derivative bosonic Lagrangian in the NR limit. In particular, we have explicitly demonstrated how to compute the contributions arising from the Metsaev–Tseytlin Lagrangian. At present, however, no mechanism is known to cancel the remaining higher-derivative divergences. We expect that the covariant rewriting presented here will provide a valuable starting point for future investigations aimed at resolving these divergences in the bosonic $\alpha^{\prime}$-corrections, as well as for extending the analysis to heterotic supergravity and more general non-relativistic gravity models.
Acknowledgements.
This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.

### v2

@@
 The construction is purely geometrical and is based on a torsionless connection.
-The non-metricities are associated with the gravitational fields of the theory, $\tau_{\mu\nu},h_{\mu\nu},\tau^{\mu\nu}$ and $h^{\mu\nu}$, and are fixed by requiring compatibility with the relativistic metric.
+The non-metricities are associated with the gravitational fields of the theory ($\tau_{\mu\nu}$, $h_{\mu\nu}$, $\tau^{\mu\nu}$, $h^{\mu\nu}$) and are fixed by requiring compatibility with the relativistic metric.
 We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature.
-Our results establish an equivalence between the proposed construction and the intrinsic torsion framework of string Newton–Cartan geometry.
+Our results establish an equivalence between the intrinsic torsion framework of string Newton–Cartan geometry and the proposed construction and.
 We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
@@
 $\tau_{\mu\nu}$ $=$ $\tau_{\mu}{}^{a}\tau_{\nu}{}^{b}\eta_{ab}\,$, (5)
-$\tau^{\mu\nu}$ $=$ $\tau^{\mu}{}_{a}\tau^{\nu}{}_{b}\eta_{ab}\,$, (6)
+$\tau^{\mu\nu}$ $=$ $\tau^{\mu}{}_{a}\tau^{\nu}{}_{b}\eta^{ab}\,$, (6)
 $h_{\mu\nu}$ $=$ $e_{\mu}{}^{a^{\prime}}e_{\nu}{}^{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (7)
-$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (8)
+$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta^{a^{\prime}b^{\prime}}\,$, (8)
 according to the vielbein decomposition of the fields ($a^{\prime}=2,...,25$). Moreover, they satisfy the defining relations,
 $\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,$,
-$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
+$\tau_{\mu}{}^{a}\tau^{\mu}{}_{b}$ $=\delta^{a}_{b}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
 Although individual geometric quantities constructed from the relativistic Levi–Civita connection exhibit divergences in the limit $c\to\infty$, the bosonic NS–NS action,
@@
 Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity.
-We expect this framework to serve as a useful starting point for higher-derivative extensions and for exploring duality-covariant formulations of string theory in the NR limit.
+We expect this framework to serve as a useful starting point for higher-derivative extensions in both supergravity and duality-covariant frameworks, in the NR limit of string theory.
 II Infinitesimal diffeomorphism transformations and covariant derivatives
@@
 $\nabla_{\mu}h_{\nu\rho}$ $=$ $Q^{(h)}_{\mu\nu\rho}\,$, (21)
-$\nabla_{\mu}h^{\nu\rho}$ $=$ $Q^{h^{-1}}_{\mu}{}^{\nu\rho}\,$. (22)
+$\nabla_{\mu}h^{\nu\rho}$ $=$ $Q^{(h^{-1})}_{\mu}{}^{\nu\rho}\,$. (22)
 The fixed values of the non-metricities come from the metric conditions,
-$\hat{\nabla}_{\mu}g_{\nu\rho}$ $=$ $0\,$,
-$\hat{\nabla}_{\mu}g^{\nu\rho}$ $=$ $0\,$,
+$\hat{\nabla}_{\mu}\hat{g}_{\nu\rho}$ $=$ $0\,$,
+$\hat{\nabla}_{\mu}\hat{g}^{\nu\rho}$ $=$ $0\,$,
 where $\hat{\nabla}$ is the usual covariant derivative with respect the relativistic Levi–Civita connection.
@@
 $Q^{(h)}_{\mu\nu\rho}$ $=$ $\tau^{\alpha}{}_{b}(\partial_{\mu}h_{(\nu\alpha}+\partial_{(\nu}h_{\mu\alpha}-\partial_{\alpha}h_{\mu(\nu})\tau_{\rho)}{}^{b}$ (25)
-$Q^{h^{-1}}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
+$Q^{(h^{-1})}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (26)
 The previous quantities are not boost invariants  [^1] Here $\lambda_{aa^{\prime}}$ is a generic boost parameter NSNS .,
@@
 The relativistic Riemann tensor is given by
-$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{(4)}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{(2)}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
-$+c^{(-2)}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{(-4)}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
+$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $c^{4}\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}+c^{2}\hat{R}^{(2)\rho}{}_{\epsilon\mu\nu}+\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ (30)
+$+c^{-2}\hat{R}^{(-2)\rho}{}_{\epsilon\mu\nu}+c^{-4}\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}\,$.
 So far we have just identified part of the covariant contribution inside $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$, given by $R^{\rho}{}_{\epsilon\mu\nu}$ but other covariant contributions are encoded in $\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ due to $c^{2}$- and $c^{-2}$-contributions in the relativistic Levi-Civita connection.
@@
 Let us start with the higher-order contribution,
-$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\nu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
+$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\partial_{\alpha}\tau_{[\mu\sigma}-\partial_{\sigma}\tau_{[\mu\alpha})$ (31)
 $(\partial_{\nu]}\tau_{\epsilon\beta}+\partial_{\epsilon}\tau_{\nu]\beta}-\partial_{\beta}\tau_{\nu]\epsilon})\,.$
 The previous quantity transforms as a tensor, and it can be written in a manifest covariant form using the covariant derivatives,
-$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
+$\hat{R}^{(4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}h^{\rho\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})$ (32)
 $(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
 The same technique can be implemented for the lowest contribution of the Riemann tensor,
-$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\nu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
+$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\partial_{\alpha}h_{[\mu\sigma}-\partial_{\sigma}h_{[\mu\alpha})$ (33)
 $(\partial_{\nu]}h_{\epsilon\beta}+\partial_{\epsilon}h_{\nu]\beta}-\partial_{\beta}h_{\nu]\epsilon})$
 which therefore can be written as
-$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
+$\hat{R}^{(-4)\rho}{}_{\epsilon\mu\nu}$ $=$ $\frac{1}{2}\tau^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})$ (34)
 $(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$
@@
 Finally, we proceed to construct the $c^{0}$-contributions to the relativistic Riemann tensor. This quantity is given by,
-$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
-$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
+$\hat{R}^{(0)\rho}{}_{\epsilon\mu\nu}$ $=R^{\rho}{}_{\epsilon\mu\nu}+\frac{1}{2}h^{\rho\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (37)
+$\qquad\quad\qquad+\frac{1}{2}\tau^{\rho\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,.$
 In the next section we will study the decomposition of the (relativistic) Ricci tensor in terms of the Newton-Cartan fields, and in covariant form with respect to infinitesimal diffeomorphisms.
@@
 Using the results of the previous subsection, we can directly write all the contributions in covariant form,
-$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\nu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
+$\hat{R}^{(4)}_{\epsilon\nu}$ $=$ $\frac{1}{2}h^{\mu\sigma}h^{\alpha\beta}(\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\alpha})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})\,$, (39)
 $\hat{R}^{(2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[h^{\mu\alpha}(\nabla_{\nu]}\tau_{\epsilon\alpha}+\nabla_{\epsilon}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\epsilon})\Big]\,$, (40)
-$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
-$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
+$\hat{R}^{(0)}_{\epsilon\nu}$= $R_{\epsilon\nu}(\tau,h)+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}\nabla_{\alpha}\tau_{[\mu\sigma}(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})$ (41)
+$\qquad\qquad+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}\nabla_{\alpha}h_{[\mu\sigma}(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})$
 $\hat{R}^{(-2)}_{\epsilon\nu}=2\nabla_{[\mu}\Big[\tau^{\mu\alpha}(\nabla_{\nu]}h_{\epsilon\alpha}+\nabla_{\epsilon}h_{\nu]\alpha}-\nabla_{\alpha}h_{\nu]\epsilon})\Big]\,.$ (42)
-$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\nu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
+$\hat{R}^{(-4)}{}_{\epsilon\nu}$ $=$ $\frac{1}{2}\tau^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\alpha})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})\,.$ (43)
 IV.3 The Ricci scalar
@@
 This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the $\hat{H}-$terms, and using the results of the previous section it can be written in the following way,
-$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+\frac{1}{2}h^{\mu\sigma}\tau^{\alpha\beta}(\nabla_{[\mu}\tau_{\alpha\sigma}+\nabla_{\alpha}\tau_{[\mu\sigma}-\nabla_{\sigma}\tau_{[\mu\sigma})(\nabla_{\nu]}h_{\epsilon\beta}+\nabla_{\epsilon}h_{\nu]\beta}-\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
-$+\frac{1}{2}\tau^{\mu\sigma}h^{\alpha\beta}(\nabla_{[\mu}h_{\alpha\sigma}+\nabla_{\alpha}h_{[\mu\sigma}-\nabla_{\sigma}h_{[\mu\sigma})(\nabla_{\nu]}\tau_{\epsilon\beta}+\nabla_{\epsilon}\tau_{\nu]\beta}-\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}$
-$+2\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\nu]}\tau_{\mu\alpha}+\nabla_{\mu}\tau_{\nu]\alpha}-\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi$
-$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
+$e^{-1}e^{2\varphi}L^{(0)}|_{NR}$ $=$ $R_{\mu\nu}(\tau,h)h^{\mu\nu}+h^{\mu\sigma}\tau^{\alpha\beta}\nabla_{\alpha}\tau_{[\mu\sigma}(\nabla_{\nu]}h_{\epsilon\beta}-\frac{1}{2}\nabla_{\beta}h_{\nu]\epsilon})h^{\epsilon\nu}$ (49)
+$+\tau^{\mu\sigma}h^{\alpha\beta}\nabla_{\alpha}h_{[\mu\sigma}(\nabla_{\epsilon}\tau_{\nu]\beta}-\frac{1}{2}\nabla_{\beta}\tau_{\nu]\epsilon})h^{\epsilon\nu}+4\nabla_{[\beta}\big[h^{\beta\alpha}(\nabla_{\mu}\tau_{\nu]\alpha}-\frac{1}{2}\nabla_{\alpha}\tau_{\nu]\mu}\big]\tau^{\mu\nu}$
+$+4h^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{2}h_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$
+$-\frac{1}{4}\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\tau^{\mu\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}\,$,
 where
@@
 The method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
-$S_{MT}=\int d^{D}x\sqrt{-g}e^{-2\phi}(L^{(0)}+L^{(1)}_{MT})\,$, (50)
+$S_{MT}=\int d^{26}x\sqrt{-g}e^{-2\phi}(L^{(0)}+L^{(1)}_{MT})\,$, (50)
 where
@@
 $+\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\rho\delta}\tau^{\sigma\lambda}h^{\nu\epsilon}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(2)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}h^{\nu\epsilon}$
-$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}$ (53)
+$+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}\,.$ (53)
 Similarly, one can inspect the rest of the terms. For $\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}$ the finite contributions are given by
@@
 $+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\gamma}\tau^{\delta\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}+8\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\lambda}h^{\epsilon\chi}h^{\alpha\psi}$
-$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}$ (56)
+$+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}\,.$ (56)
 At the present moment we cannot control all the divergences of the full Metsaev-Tseytlin Lagrangian and for that reason we are not giving an exhaustive analysis of the four-derivative contributions of this Lagrangian in its NR limit.

### v3

@@
 Abstract
-We construct a diffeomorphism-covariant formulation of the non-relativistic (NR) limit of bosonic supergravity.
-This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to the NR degrees of freedom.
-The construction is purely geometrical and is based on a torsionless connection.
-The non-metricities are associated with the gravitational fields of the theory ($\tau_{\mu\nu}$, $h_{\mu\nu}$, $\tau^{\mu\nu}$, $h^{\mu\nu}$) and are fixed by requiring compatibility with the relativistic metric.
+We construct a metric-like formulation of the non-relativistic (NR) limit of bosonic supergravity at the Lagrangian level.
+This formulation is particularly useful for decomposing relativistic tensors, such as powers of the Riemann tensor, in a manifest covariant form with respect to infinitesimal diffeomorphisms.
+The construction is purely geometrical and is based on a torsionless connection, mimicking the construction of the relativistic theory.
+The formulation contains non-vanishing non-metricities, which are associated with the gravitational fields of the theory ($\tau_{\mu\nu}$, $h_{\mu\nu}$, $\tau^{\mu\nu}$, $h^{\mu\nu}$).
+The non-metricities are fixed by requiring compatibility with the relativistic metric, before taking the NR expansion.
 We provide a fully covariant decomposition of the relativistic Riemann tensor, Ricci tensor, and scalar curvature.
-Our results establish an equivalence between the intrinsic torsion framework of string Newton–Cartan geometry and the proposed construction and.
-We also discuss potential applications, including a manifestly diffeomorphism-covariant rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving bosonic $\alpha^{\prime}$-corrections under the same limit, and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
+Our results establish an equivalence between the vielbein approach of string Newton–Cartan geometry at the level of the Lagrangian and the proposed construction.
+We also discuss potential applications, including a pure metric rewriting of the two-derivative finite bosonic supergravity Lagrangian under the NR limit, a powerful simplification in deriving NR bosonic $\alpha^{\prime}$-corrections and extensions to more general $f(R,Q)$ Newton–Cartan geometries.
 I Introduction
-In recent years, non-relativistic (NR) limits of string theory NR1 -NR3 and supergravity NRST1 -NRST13 have attracted considerable attention, motivated by both conceptual and more phenomenological developments.
-In particular, the systematic understanding of the supergravity limit of NR string theory has led to the emergence of Newton–Cartan and string Newton–Cartan geometries as the natural geometric frameworks underlying these limits (for reviews and complementary introductions, see Review -Review3 and references therein).
-Let us start by considering the universal bosonic NS–NS supergravity, whose relativistic field content consists of a spacetime metric $\hat{g}_{\mu\nu}$, a Kalb–Ramond two-form $\hat{B}_{\mu\nu}$ and a dilaton $\hat{\phi}$.
-A controlled NR limit is obtained NSNS by introducing an explicit expansion in powers of the speed of light $c$,
+In recent years, non-relativistic (NR) limits of string theory NR1 -NR3 and supergravity NRST1 -NRST13 have attracted considerable attention, motivated by both conceptual developments and potential phenomenological applications.
+In particular, the systematic study of the supergravity limit of NR string theory has led to the emergence of Newton–Cartan and string Newton–Cartan geometries as the natural geometric frameworks underlying these limits (for reviews and complementary introductions, see Review -Review3 and references therein).
+While the NR limit of NS-NS gravity has been deeply studied in NSNS , the authors have addressed the construction in the vielbein formalism of NR theory.
+In this work we will reformulate this same setup, but using a fully metric approach, where the full Lorentz symmetry is manifest from the starting point.
+In other words, we will address NSNS-gravity in its metric formulation directly form the relativistic theory, and then we will study extensions which keep this structure.
+This way to avoid the vielbein formulation will be extremely useful to address higher-derivative contributions in its NR limit, which can be written entirely with curvatures, like powers of the Riemann tensor.
+In this sense, one can avoid the use of the spin connection, and work solely with a affine connection.
+Our canonical example here is the Metsaev and Tseytlin formulation of the four-derivative corrections of bosonic string theory MetsaevTseytlin .
+Let us begin by considering the universal bosonic NS–NS supergravity, whose relativistic field content consists of a spacetime metric $\hat{g}_{\mu\nu}$, a Kalb–Ramond two-form $\hat{B}_{\mu\nu}$ and a dilaton $\hat{\phi}$.
+A controlled NR limit can be obtained NSNS by introducing an explicit expansion in powers of $c$,
 $\hat{g}_{\mu\nu}$ $=$ $c^{2}\tau_{\mu\nu}+h_{\mu\nu}\,$, (1)
 $\hat{g}^{\mu\nu}$ $=$ $\frac{1}{c^{2}}\tau^{\mu\nu}+h^{\mu\nu}\,$, (2)
-$\hat{B}_{\mu\nu}$ $=$ $-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,$, (3)
+$\hat{B}_{\mu\nu}$ $=$ $-c^{2}c_{\mu\nu}+b_{\mu\nu}\,$, (3)
 $\hat{\phi}$ $=$ $\ln(c)+\varphi\,$, (4)
-where $\mu,\nu=0,\dots,25$ are space-time indices and the longitudinal directions are labeled by $a,b=0,1$. We also have
-$\tau_{\mu\nu}$ $=$ $\tau_{\mu}{}^{a}\tau_{\nu}{}^{b}\eta_{ab}\,$, (5)
-$\tau^{\mu\nu}$ $=$ $\tau^{\mu}{}_{a}\tau^{\nu}{}_{b}\eta^{ab}\,$, (6)
-$h_{\mu\nu}$ $=$ $e_{\mu}{}^{a^{\prime}}e_{\nu}{}^{b^{\prime}}\eta_{a^{\prime}b^{\prime}}\,$, (7)
-$h^{\mu\nu}$ $=$ $e^{\mu}{}_{a^{\prime}}e^{\nu}{}_{b^{\prime}}\eta^{a^{\prime}b^{\prime}}\,$, (8)
-according to the vielbein decomposition of the fields ($a^{\prime}=2,...,25$). Moreover, they satisfy the defining relations,
-$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,$,
-$\tau_{\mu}{}^{a}\tau^{\mu}{}_{b}$ $=\delta^{a}_{b}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (9)
-Although individual geometric quantities constructed from the relativistic Levi–Civita connection exhibit divergences in the limit $c\to\infty$, the bosonic NS–NS action,
+where $\mu,\nu=0,\dots,25$ are space-time indices and $c_{\mu\nu}$ satisfies $c_{\mu\rho}\tau^{\rho\sigma}c_{\sigma\nu}=\tau_{\mu\nu}$. These variables satisfy the defining relations
+$h^{\mu\nu}\tau_{\nu\rho}$ $=$ $h^{\mu\nu}c_{\nu\rho}=0\,$, (5)
+$h_{\mu\nu}\tau^{\nu\rho}$ $=$ $0\,$, (6)
+$h_{\mu\nu}h^{\nu\rho}$ $=$ $\delta_{\mu}^{\rho}-\tau_{\mu\nu}\tau^{\nu\rho}\,$, (7)
+Although individual geometric quantities constructed from the relativistic Levi–Civita connection diverge in the limit $c\to\infty$, the bosonic NS–NS action
 $S=\int d^{26}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big)\,$, (8)
-with $\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}$, remains finite due to a non-trivial cancellation between divergent contributions arising from the Ricci scalar and the Kalb–Ramond sector.
-This observation has been central to the development of other NR string supergravity theories such as heterotic formulations BandR - Sigmaheterotic , supersymmetric extensions NRST11 -NRST12 and Double Field Theory DFT1 -DFT4 , in the context of non-Riemannian double geometries NRDFT1 -NRDFT9 and its NR limit EandD .
-Motivated by the progress on the NR limit of string supergravities, different approaches to study higher-derivative corrections under the NR limit have recently emerged Higher-Derivative1 -Higher-Derivative4 .
-Some of the theories already explored can be written in terms of tensors with respect to infinitesimal diffeomorphisms in their relativistic level.
-In order to be able to decompose these tensors preserving the covariance, one needs to construct a suitable connection with respect to the longitudinal and transverse parts of the metric degrees of freedom (and its inverse), and rewrite these contributions in terms of curvatures and covariant derivatives with respect to this symmetry.
-The same can be done for the Lorentz inherited symmetries, like boosts, $SO(1,1)$ and $SO(8)$ transformations, which was done for the NR limit of bosonic supergravity in NSNS at the two-derivative level.
-In particular, a fully diffeomorphism-covariant formulation of this theory, formulated directly in terms of intrinsic NR geometric data and their associated curvatures is still lacking.
-The main obstruction lies in the fact that the relativistic Levi–Civita connection does not admit a smooth NR limit when expressed solely in terms of Newton–Cartan variables.
-In this work, we solve this issue by constructing a torsionless affine connection that is explicitly covariant under spacetime diffeomorphisms.
-Crucially, this connection is not metric compatible with respect to the fundamental NR fields.
-Instead, the covariant derivatives of $\tau_{\mu}{}^{a}$ and $h^{\mu\nu}$ are controlled by specific non-metricity tensors, which are uniquely fixed by requiring consistency with the relativistic metric condition.
-As a result, the full bosonic supergravity action can be rewritten in a manifestly covariant form with respect to diffeomorphisms, where all terms remain finite and well defined as $c\rightarrow\infty$.
+with $\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}$, remains finite due to a non-trivial cancellation between divergent contributions arising from the Ricci scalar and the Kalb–Ramond sector (in vielbein formalims $c_{\mu\nu}=\tau_{\mu}{}^{a}\tau_{\nu}{}^{b}\epsilon_{ab}$ ensures the cancellation, with $a,b=0,1$ the transversal flat directions).
+This mechanism underlies the construction of various non-relativistic string supergravity theories, including heterotic formulations BandR -Sigmaheterotic , supersymmetric extensions NRST11 -NRST12 and Double Field Theory constructions DFT1 -DFT4 , as well as their formulation within non-Riemannian double geometries NRDFT1 -NRDFT9 and their NR limits EandD .
+Motivated by these developments, increasing attention has recently been devoted to understanding higher-derivative corrections in the NR limit Higher-Derivative1 -Higher-Derivative4 .
+At the relativistic level, many of these corrections are naturally written in terms of covariant tensors constructed from the Levi–Civita connection.
+In order to study their NR limit while preserving covariance under spacetime diffeomorphisms, it is useful to express these objects in terms of curvatures and covariant derivatives adapted to the NR geometric variables appearing in the expansion (1).
+Previous formulations of the NR limit of NS–NS gravity provide a consistent description of the dynamics in terms of Newton–Cartan geometry and connections that naturally accommodate the local boost, $SO(1,1)$ and transverse rotation symmetries of the theory NSNS .
+While this framework is well suited for describing the two-derivative dynamics of NR supergravity certain applications, in particular the systematic decomposition of relativistic curvature tensors appearing in higher-derivative corrections, benefit from a pure metric formulation.
+In such a formulation, the relativistic curvature invariants could be reorganized in a covariant manner in terms of their different contributions in powers of $c$ without need of using the vielbein decomposition or spin connections (see also new for a metric formalism in p-brane Galilean geometries).
+The main goal of this work is to provide such a formulation.
+We construct a torsionless affine connection adapted to the NR metric variables $\tau_{\mu\nu}$ and $h_{\mu\nu}$ and use it to rewrite relativistic geometric quantities in a manifestly covariant form under spacetime diffeomorphisms.
+Importantly, the resulting connection is not metric compatible with respect to the fundamental NR fields.
+Instead, the covariant of the fundamental fields are controlled by a set of non-metricity tensors that arise naturally from expanding the relativistic metric compatibility condition.
+These non-metricities are uniquely determined by requiring consistency with the relativistic Levi–Civita structure (compatibility with the relativistic metric).
+This construction allows relativistic curvature tensors to be systematically reorganized into NR geometric invariants.
+In particular, we show that the full bosonic supergravity action can be rewritten in terms of covariant NR quantities, with all contributions remaining finite and well defined in the limit $c\rightarrow\infty$.
 The purpose of this paper is threefold.
-First, we present an explicit construction of a torsionless connection and the associated non-metricities within the framework of Newton–Cartan gravity.
-Second, we compute the corresponding curvatures and show explicitly how all relativistic curvature tensors reorganize into NR geometric invariants.
-Third, we discuss several applications of the formalism.
-These include the derivation of the complete non-relativistic bosonic supergravity Lagrangian in a fully covariant form, the systematic analysis of higher-derivative contributions such as Riemann-squared and other four-derivative bosonic corrections, and the construction of more general Newton–Cartan–inspired theories with alternative choices of non-metricity that relax boost invariance beyond the bosonic supergravity setting.
-Our results provide a geometric foundation for non-relativistic bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally accommodating the presence of non-metricity.
-We expect this framework to serve as a useful starting point for higher-derivative extensions in both supergravity and duality-covariant frameworks, in the NR limit of string theory.
-II Infinitesimal diffeomorphism transformations and covariant derivatives
+First, we present an explicit construction of a torsionless affine connection and the associated non-metricity tensors within the Newton–Cartan framework.
+Second, we compute the corresponding curvature tensors and demonstrate how relativistic curvature invariants decompose into NR geometric objects in a covariant manner.
+Third, we illustrate several applications of the formalism.
+These include the derivation of the complete NR two-derivative bosonic supergravity Lagrangian in a manifestly covariant form, the analysis of the finite form of the Metsaev-Tseytlin alpha’-contributions, and the construction of more general Newton–Cartan-inspired theories based on alternative choices of non-metricity that extend beyond the specific structure inherited from bosonic supergravity.
+Overall, our results provide a geometric framework for the NR limit of bosonic supergravity that parallels the role played by the Levi–Civita connection in relativistic theories, while naturally incorporating non-metricity effects.
+We expect this formalism to provide a useful starting point for the systematic study of higher-derivative corrections and duality-covariant formulations in the non-relativistic limit of string theory.
+II Pure metric formalism and the affine connection
 We start by recalling the transformation rule of the fundamental fields with respect to infinitesimal diffeomorphisms,
-$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau_{\mu}{}^{a}}+\partial_{\mu}{\xi^{\rho}}\tau_{\rho}{}^{a}\,$,
-$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau^{\mu}{}_{a}}-\partial_{\rho}{\xi^{\mu}}\tau^{\rho}{}_{a}\,$,
-$\delta_{\xi}h_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h_{\mu\nu}}+2\partial_{(\mu}{\xi^{\rho}}h_{\nu)\rho}\,$,
+$\delta_{\xi}\tau_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau_{\mu}{}^{a}}+2\partial_{(\mu|}{\xi^{\rho}}\tau_{\rho|\nu)}\,$,
+$\delta_{\xi}\tau^{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}\tau^{\mu\nu}-2\partial_{\rho}{\xi^{(\mu}}\tau^{\rho\nu)}\,$,
+$\delta_{\xi}h_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}h_{\mu\nu}+2\partial_{(\mu|}\xi^{\rho}h_{|\nu)\rho}\,$,
 $\delta_{\xi}h^{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h^{\mu\nu}}–2\partial_{\rho}{\xi^{(\mu}}h^{\nu)\rho}\,.$ (9)
@@
 $\Gamma_{\mu\nu}^{\rho}(\tau,h)=\frac{1}{2}h^{\rho\sigma}(2\partial_{(\mu}h_{\nu)\sigma}-\partial_{\sigma}h_{\mu\nu})$
-[…]
+$+\frac{1}{2}\tau^{\rho\sigma}(2\partial_{(\mu}(\tau_{\nu)\sigma}-\partial_{\sigma}(\tau_{\mu\nu}))\,$, (10)
 which is the $c^{0}$-contribution coming from the relativistic connection, preserving the non-tensorial transformation rule, in the sense that it is the contribution which transforms as a connection.
 The covariant derivatives are therefore defined as
-$\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}\,$, (13)
-$\nabla_{\mu}\tau^{\nu}{}_{a}$ $=$ $\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\sigma}^{\nu}\tau^{\sigma}{}_{a}\,$, (14)
+$\nabla_{\mu}\tau_{\nu\sigma}$ $=$ $\partial_{\mu}\tau_{\nu\sigma}-2\Gamma_{\mu(\nu|}^{\rho}\tau_{\rho|\sigma)}\,$, (11)
+$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $\partial_{\mu}\tau^{\nu\rho}+2\Gamma_{\mu\sigma}^{(\nu|}\tau^{\sigma|\rho)}\,$, (12)
 $\nabla_{\mu}h_{\nu\rho}$ $=$ $\partial_{\mu}h_{\nu\rho}-2\Gamma_{\mu(\nu}^{\sigma}h_{\rho)\sigma}\,$, (13)
 $\nabla_{\mu}h^{\nu\rho}$ $=$ $\partial_{\mu}h^{\nu\rho}+2\Gamma_{\mu\sigma}^{(\nu}h^{\rho)\sigma}\,.$ (14)
-The Riemann tensor can be easily computed from the commutator of the covariant derivatives,
-$[\nabla_{\mu},\nabla_{\nu}]\tau^{\rho}{}_{a}=R^{\rho}{}_{\epsilon\mu\nu}\tau^{\epsilon}{}_{a}\,$, (17)
+The Riemann tensor can be easily computed from the commutator of the covariant derivatives acting on an arbitrary vector $v^{\mu}$,
+$[\nabla_{\mu},\nabla_{\nu}]v^{\rho}=R^{\rho}{}_{\epsilon\mu\nu}v^{\epsilon}\,$, (15)
 giving the usual expression
@@
 $Q^{(h^{-1})}_{\rho}{}^{\mu\nu}$ $=$ $-h^{(\mu\alpha}(2\partial_{(\rho}\tau_{\sigma)\alpha}-\partial_{\alpha}\tau_{\rho\sigma})\tau^{\sigma\nu)}\,.$ (24)
-The previous quantities are not boost invariants  [^1] Here $\lambda_{aa^{\prime}}$ is a generic boost parameter NSNS .,
-$\delta_{\lambda}h_{\mu\nu}$ $=$ $-2\lambda_{aa^{\prime}}e_{(\mu}{}^{a^{\prime}}\tau_{\nu)}{}^{a}\,$, (27)
-$\delta_{\lambda}\tau^{\mu\nu}$ $=$ $2\lambda_{a}{}^{a^{\prime}}e^{(\mu}{}_{a^{\prime}}\tau^{\nu)a}\,$, (28)
-$\delta_{\lambda}b_{\mu\nu}$ $=$ $-2\epsilon_{ab}\lambda^{a}{}_{a^{\prime}}\tau_{[\mu}{}^{b}e_{\nu]}{}^{a^{\prime}}\,$. (29)
-and consequently the non-metricities cannot be set to zero without compromising the boost symmetry in bosonic supergravity.
+The previous quantities are not boost invariants and consequently the non-metricities cannot be set to zero without compromising the boost symmetry in bosonic supergravity.
 We will return to this point in the discussion section, since one might be interested in exploring Newton-Cartan geometries with more arbitrary non-metricities, beyond bosonic supergravity.
@@
 This choice explicitly breaks the boost symmetry of the NR limit of bosonic supergravity.
-V.1 NR-Bosonic supergravity Lagrangian: covariant (re-)formulation
+V.1 NR-Bosonic supergravity Lagrangian: metric-like (re-)formulation
 This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the $\hat{H}-$terms, and using the results of the previous section it can be written in the following way,
@@
 where
-$h_{\mu\nu\rho}$ $=$ $3\partial_{[\mu}b_{\nu\rho]}\,,\quad\hat{H}^{(2)}_{\mu\nu\rho}=-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,$.
+$h_{\mu\nu\rho}$ $=$ $3\partial_{[\mu}b_{\nu\rho]}\,,\quad\hat{H}^{(2)}_{\mu\nu\rho}=-3\nabla_{[\mu}c_{\nu\rho]}\,.$
 Now we can explicitly see that the finite bosonic Lagrangian after taking the NR limit contains a $R_{\mu\nu}h^{\mu\nu}$ contribution, plus other terms related to the non-metricities.
-The construction is equivalent to the torsionful-formalism of string Newton-Cartan.
+As we mentioned before, the construction is purely written in metric formalism and is equivalent to the torsionful-formalism of string Newton-Cartan.
+It is worth-mentioning that the compatibility explored in this paper matches at the level of the action.
+The compatibility at the equations of motion is not guaranteed, since it is not clear that the torsion constraint found in NSNS is required in the metric approach.
 V.2 Non-relativistic bosonic $\alpha^{\prime}$-corrections
@@
 $+4\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}h_{\mu\gamma}h^{\sigma\lambda}+2\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau_{\mu\gamma}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\hat{R}^{(4)\gamma}{}_{\epsilon\delta\lambda}\tau^{\nu\epsilon}\tau^{\rho\delta}\tau^{\sigma\lambda}h_{\mu\gamma}\,.$ (48)
-Similarly, one can inspect the rest of the terms. For $\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}$ the finite contributions are given by
-$\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\gamma\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\sigma\alpha}h_{\mu\gamma\delta}h_{\epsilon\lambda\alpha}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\rho\gamma}h^{\nu\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\epsilon\lambda}h_{\gamma\delta\alpha}$
-$+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\sigma}{}_{\gamma\epsilon\delta}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}h^{\nu\epsilon}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
-$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\gamma\lambda}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\lambda}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
-$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\gamma\lambda}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}$
-$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\mu}{}_{\sigma\gamma\epsilon}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}$
-$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}\tau^{\gamma\lambda}h^{\rho\sigma}h^{\epsilon\alpha}h_{\delta\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}\tau^{\epsilon\lambda}h^{\nu\sigma}h^{\rho\alpha}h_{\delta\lambda\alpha}$
-$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\gamma\delta}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(-2)\mu}{}_{\delta\lambda\alpha}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
-$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}$
-$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}h^{\rho\delta}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
-$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}h^{\rho\gamma}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}\tau^{\epsilon\alpha}h^{\rho\delta}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}\tau^{\gamma\lambda}h^{\epsilon\alpha}$
-$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\rho\epsilon}\,.$ (54)
+Due to the lenght of the expressions, we provide the remaining H-dependent terms in the Appendix.
+V.3 General f(R,Q) non-relativistic theories
+While in this work we have shown that the non-metricities (21)-(24) need to be fixed in a very particular way to provide the relativistic metric conditions (III)-(III), one can construct more general NR theories with arbitrary non-metricities.
+One interesting example is to demand the compatibility condition on the Newton-Cartan fields,
+$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $0\,$, (49)
+$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $0\,$, (50)
+$\nabla_{\mu}h_{\nu\rho}$ $=$ $0\,$, (51)
+$\nabla_{\mu}h^{\nu\rho}$ $=$ $0\,$, (52)
+which automatically implies that both the relativistic Riemann and Ricci tensor are now finite,
+$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $R^{\rho}{}_{\epsilon\mu\nu}\,,\quad\hat{R}_{\epsilon\nu}=R_{\epsilon\nu}\,$, (53)
+and the more general (gravitational) Lagrangian in this setup is given by
+$L_{NR}|_{Q=0}=a_{1}R_{\mu\nu}h^{\mu\nu}+a_{2}R_{\mu\nu}\tau^{\mu\nu}\,$, (54)
+with $a_{1}$ and $a_{2}$ two arbitrary coefficients.
+In this case, the $SO(8)$ and $SO(1,1)$ symmetry from the vielbein formalism can be preserved, but the boost symmetry is broken.
+Further theories with these properties can be explored by changing the non-metricity contributions and including $L(R,Q)$ in the $L_{NR}$ Lagrangian.
+VI Discussion
+The present construction provides a geometrical formulation of the non-relativistic limit of bosonic supergravity in terms of a Newton–Cartan geometry with non-metricities.
+Within this framework, an appropriate choice of connection allows a metric formulation, mimicking the construction of the relativistic cases.
+As a result, the full two-derivative bosonic supergravity in its NR limit can be written in a covariant form with respect to infinitesimal diffeomorphisms, thereby complementing other approaches based on subgroups of the Lorentz symmetry.
+Moreover, we find compatibility of the approaches at the level of the action, while the study of the equations of motion deserve further study to understand if it is necessary a new constraint, equivalent to the torsion constraint NSNS .
+A key advantage of the construction is that it allows for a straightforward decomposition of bosonic higher-derivative contributions, such as $\alpha^{\prime}$-corrections.
+These corrections are naturally formulated in the metric formalism, where the Lorentz symmetry is manifest and no vielbein (or spin connection) is required.
+From a practical perspective, this provides a systematic method for analyzing higher-derivative terms by expressing relativistic curvature invariants directly in terms of non-relativistic geometric data.
+In particular, the use of fixed non-metricities enables a transparent identification of divergent and finite contributions, a task that is typically challenging.
+Although we do not yet have full control over the divergences that appear at the four-derivative level, the formalism developed in this work already proves useful for isolating and organizing several finite contributions to the four-derivative bosonic Lagrangian in the NR limit.
+In particular, we have explicitly demonstrated how to compute the contributions arising from the Metsaev–Tseytlin Lagrangian.
+At present, however, no mechanism is known to cancel the remaining higher-derivative divergences.
+We expect that the covariant rewriting presented here will provide a valuable starting point for future investigations aimed at resolving these divergences in the bosonic $\alpha^{\prime}$-corrections, as well as for extending the analysis to heterotic supergravity and more general non-relativistic gravity models.
+Acknowledgements.
+This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
+The author is very grateful to Jan Rosseel for discussions and clarifications on the article NSNS .
+Appendix A Full finite Metsaev and Tseytlin Lagrangian-H-dependent terms
 The terms only containing $\hat{H}-$contributions can also be easily written in covariant form. For example, $H^{4}$ contributes with the following finite terms,
@@
 $+4\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{H}^{(2)}_{\delta\lambda\alpha}\hat{H}^{(2)}_{\beta\chi\psi}\tau^{\mu\sigma}\tau^{\nu\delta}\tau^{\gamma\beta}h^{\rho\epsilon}h^{\lambda\chi}h^{\alpha\psi}\,.$ (56)
+For $\hat{H}^{\mu\nu\rho}\hat{H}_{\mu\sigma\lambda}\hat{R}_{\nu\rho}{}^{\sigma\lambda}$ the finite contributions are given by
+$\hat{R}^{(2)\mu}{}_{\nu\rho\sigma}h^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\gamma\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\sigma\alpha}h_{\mu\gamma\delta}h_{\epsilon\lambda\alpha}+2\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\rho\gamma}h^{\nu\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\epsilon\lambda}h_{\gamma\delta\alpha}$
+$+\hat{R}^{(4)\mu}{}_{\nu\rho\sigma}\tau^{\nu\gamma}h^{\rho\epsilon}h^{\sigma\delta}h^{\lambda\alpha}h_{\mu\gamma\lambda}h_{\epsilon\delta\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\sigma}{}_{\gamma\epsilon\delta}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}h^{\nu\epsilon}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
+$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}h^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\sigma}{}_{\gamma\epsilon\delta}\tau^{\gamma\lambda}h^{\mu\epsilon}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\lambda}h^{\rho\delta}h^{\gamma\alpha}h_{\sigma\lambda\alpha}$
+$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\nu\delta}h^{\rho\lambda}h^{\gamma\alpha}h_{\sigma\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}h_{\sigma\lambda\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\sigma}{}_{\gamma\epsilon\delta}\tau^{\mu\epsilon}\tau^{\gamma\lambda}h^{\nu\delta}h^{\rho\alpha}h_{\sigma\lambda\alpha}$
+$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(0)\mu}{}_{\sigma\gamma\epsilon}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}h^{\nu\sigma}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}$
+$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(2)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\delta}\tau^{\gamma\lambda}h^{\rho\sigma}h^{\epsilon\alpha}h_{\delta\lambda\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\gamma\delta}\tau^{\epsilon\lambda}h^{\nu\sigma}h^{\rho\alpha}h_{\delta\lambda\alpha}$
+$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{R}^{(4)\mu}{}_{\sigma\gamma\epsilon}\tau^{\nu\sigma}\tau^{\gamma\delta}h^{\rho\lambda}h^{\epsilon\alpha}h_{\delta\lambda\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(-2)\mu}{}_{\delta\lambda\alpha}h^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
+$-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}h^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}+2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}h^{\nu\gamma}h^{\rho\delta}h^{\epsilon\alpha}+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(0)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}h^{\rho\sigma}h^{\gamma\lambda}h^{\epsilon\alpha}$
+$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}h^{\rho\delta}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\nu\epsilon}h^{\rho\delta}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}h^{\gamma\lambda}h^{\epsilon\alpha}$
+$-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(2)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}h^{\rho\gamma}h^{\epsilon\alpha}-\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\gamma\lambda}\tau^{\epsilon\alpha}h^{\rho\delta}-2\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\sigma}\tau^{\rho\delta}\tau^{\gamma\lambda}h^{\epsilon\alpha}$
+$+\hat{H}^{(2)}_{\mu\nu\rho}\hat{H}^{(2)}_{\sigma\gamma\epsilon}\hat{R}^{(4)\mu}{}_{\delta\lambda\alpha}\tau^{\nu\delta}\tau^{\sigma\lambda}\tau^{\gamma\alpha}h^{\rho\epsilon}\,.$ (57)
 At the present moment we cannot control all the divergences of the full Metsaev-Tseytlin Lagrangian and for that reason we are not giving an exhaustive analysis of the four-derivative contributions of this Lagrangian in its NR limit.
 However, the covariant framework developed in this work could be very useful for attacking this open problem.
-V.3 General f(R,Q) non-relativistic theories
-While in this work we have shown that the non-metricities (23)-(26) need to be fixed in a very particular way to provide the relativistic metric conditions (III)-(III), one can construct more general NR theories with arbitrary non-metricities.
-One interesting example is to demand the compatibility condition on the Newton-Cartan fields,
-$\nabla_{\mu}\tau_{\nu\rho}$ $=$ $0\,$, (57)
-$\nabla_{\mu}\tau^{\nu\rho}$ $=$ $0\,$, (58)
-$\nabla_{\mu}h_{\nu\rho}$ $=$ $0\,$, (59)
-$\nabla_{\mu}h^{\nu\rho}$ $=$ $0\,$, (60)
-which automatically implies that both the relativistic Riemann and Ricci tensor are now finite,
-$\hat{R}^{\rho}{}_{\epsilon\mu\nu}$ $=$ $R^{\rho}{}_{\epsilon\mu\nu}\,$, (61)
-$\hat{R}_{\epsilon\nu}$ $=$ $R_{\epsilon\nu}\,$, (62)
-and the more general (gravitational) Lagrangian in this setup is given by
-$L_{NR}|_{Q=0}=a_{1}R_{\mu\nu}h^{\mu\nu}+a_{2}R_{\mu\nu}\tau^{\mu\nu}\,$, (63)
-with $a_{1}$ and $a_{2}$ two arbitrary coefficients.
-In this case, the $SO(8)$ and $SO(1,1)$ symmetry from the vielbein formalism can be preserved, but the boost symmetry is broken.
-Further theories with these properties can be explored by changing the non-metricity contributions and including $L(R,Q)$ in the $L_{NR}$ Lagrangian.
-VI Discussion
-The present construction provides a geometrical formulation of the non-relativistic limit of bosonic supergravity in terms of a Newton–Cartan geometry with non-metricities.
-Within this framework, an appropriate choice of connection allows spacetime diffeomorphism covariance to remain manifest throughout the construction.
-As a result, the full two-derivative bosonic supergravity in its NR limit can be written in a covariant form with respect to infinitesimal diffeomorphisms, thereby complementing other approaches based on subgroups of the Lorentz symmetry NSNS .
-A key advantage of this formalism is that it allows for a straightforward decomposition of bosonic higher-derivative contributions, such as $\alpha^{\prime}$-corrections.
-These corrections are naturally formulated in the metric formalism, where the Lorentz symmetry is manifest and no vielbein (or spin connection) is required.
-From a practical perspective, this provides a systematic method for analyzing higher-derivative terms by expressing relativistic curvature invariants directly in terms of non-relativistic geometric data.
-In particular, the use of fixed non-metricities enables a transparent identification of divergent and finite contributions, a task that is typically challenging.
-Although we do not yet have full control over the divergences that appear at the four-derivative level, the formalism developed in this work already proves useful for isolating and organizing several finite contributions to the four-derivative bosonic Lagrangian in the NR limit.
-In particular, we have explicitly demonstrated how to compute the contributions arising from the Metsaev–Tseytlin Lagrangian.
-At present, however, no mechanism is known to cancel the remaining higher-derivative divergences.
-We expect that the covariant rewriting presented here will provide a valuable starting point for future investigations aimed at resolving these divergences in the bosonic $\alpha^{\prime}$-corrections, as well as for extending the analysis to heterotic supergravity and more general non-relativistic gravity models.
-Acknowledgements.
-This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.